# SAC: A Framework for Measuring and Inducing Personality Traits in LLMs with Dynamic Intensity Control

### v1

@@ 8 Conclusion and Future Work
Our study demonstrates that continuous, intensity-based personality control , enabled by the SAC framework and grounded in the 16PF model, offers a robust and interpretable method for modulating LLM personas. Through systematic evaluations across neutral baselines, binary induction (P2), and graded induction (SAC), we show that LLMs not only respond predictably to personality modulation but also reflect inter-trait dynamics aligned with human psychometric theory. This establishes a strong foundation for building adaptive, psychologically informed AI agents. Future work will explore multi-trait induction to simulate more human-like, multidimensional personalities. Another promising direction is real-time dynamic modulation of personality based on user interaction, enabling context-aware adaptations during dialogue.
{ack}

### v2

@@ 8 Conclusion and Future Work
Our study demonstrates that continuous, intensity-based personality control, enabled by the SAC framework and grounded in the 16PF model, offers a robust and interpretable method for modulating LLM personas. Through systematic evaluations across neutral baselines, binary induction (P2), and graded induction (SAC), we show that LLMs not only respond predictably to personality modulation but also reflect inter-trait dynamics aligned with human psychometric theory. This establishes a strong foundation for building adaptive, psychologically informed AI agents. Future work will explore multi-trait induction to simulate more human-like, multidimensional personalities. Another direction is real-time dynamic modulation of personality based on user interaction, enabling context-aware adaptations during dialogue.
ACKNOWLEDGEMENTS